OpenStudy (anonymous):

3x+5y=-5 i want to get the y be it self

OpenStudy (anonymous):

5y = -5 -3x y = (-5 - 3x)/(5)

OpenStudy (anonymous):

i dont get it so its y=3/5x1

OpenStudy (anonymous):

\[3x + 5y = -5\] \[5y = -5 -3x\] \[y = \frac{ -5 - 3x }{ 5 }\]

OpenStudy (anonymous):

You're trying to separate y and put all the x's on the other side, So you first minus 3x on both sides, Then you divide both sides by 5 to get y alone

OpenStudy (anonymous):

You understand? o.o

OpenStudy (anonymous):

yeah thank you